A circular current carrying coil has a radius R. The distance from the centre of the coil, on the axes, where B will be 1/8 of its value at the centre of the coil is

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Magnetic field at the center of circular coil, B = μoL/2R.
Magnetic field at a distance r from the circular coil,
          B = μoR2L/2(R2 + r2)3/2
And as per the question μoR2L/2(R2 + r2)3/2  = (1/8)(μ oL/2R)
ð 8R3 = (R2 + r2)3/2
ð R2 + r2 = 4R2
ð R = √3 R